OpenStudy (anonymous):

Show the work to solve |5x – 2| > 8

OpenStudy (anonymous):

would you still like some assistance with this problem?

OpenStudy (anonymous):

Please :)

OpenStudy (anonymous):

okeedokee

OpenStudy (anonymous):

Im just not sure how to set it up..

OpenStudy (anonymous):

there are two possible linear equations that result from this inequality.

OpenStudy (anonymous):

either 5x-2>8 or 5x-2<-8.

OpenStudy (anonymous):

what the inequality says is that the distance from 5x to 2 is greater than 8 units.

OpenStudy (anonymous):

Something did not show up when i typed in the equation it was supposed to be |5x - 2| greater than or equal to 8... Im sorry.

OpenStudy (anonymous):

ok. 5x-2>or=8 or 5x-2<or=-8

OpenStudy (anonymous):

\[\ge \]

OpenStudy (anonymous):

the equation is the condition. the distance is the 8. how far 5x is from 2 is at least 8 units. nothing really changes with the equals sign in terms of set up

OpenStudy (anonymous):

solve the two above inequalities and you will obtain your solutions. there should be two

OpenStudy (anonymous):

i had one on here similar to this one and they set it up kind of strange, is that how this one is going to be set up too?

OpenStudy (anonymous):

what is the other problem?

OpenStudy (anonymous):

the direction of the inequality determines the equations. for instance, |x-5|>2 gives the two inequalities x-5>2 or x-5<-2

OpenStudy (anonymous):

whereas the inequality |x-5|<2 leads to the equations x-5>-2 and x-5<2 or the compound inequality -2<x-5<2

OpenStudy (anonymous):